Suppose you are told that a 95% confidence interval for the average price of a gallon of regular gasoline in your state is from $2.91 to $4.40. Use the fact that the confidence interval for the mean is in the form xE to x + E to compute the sample mean and the maximal margin of error E. (Round your answers to two decimal places.)

x = $
E = $
Answer #1

Solution:

Margin of error Can be calculated as

Margin of error = (Upper bound - Lower bound)/2 = (4.40-2.91)/2 = 0.745

So sample mean = (upper bound - margin of error) = 4.40-0.745 = 3.655

Sample mean = 3.66

Margin of error = 0.75